OpenStudy (anonymous):

Find the circumference and area of a circle with a diameter of 10 inches. Leave your answers in terms of pi. (4 points) C = 5π; A = 20π C = 10π; A = 20π C = 10π; A = 25π C = 5π; A = 25π

OpenStudy (jackofalltradez):

Circumference of a circle = 2(pi)r So, r = 5, 5 * 2 = 10, so the final answer for C is 10(pi) Area of a circle = (pi)r^2 So, r = 5, 5 squared = 25, so the final answer for A is 25(pi) You can figure out the rest from there :) @grex10

OpenStudy (anonymous):

so the last choice is right?

OpenStudy (anonymous):

@JackofallTradez

OpenStudy (jackofalltradez):

No, the third to last choice is right, because it is 'C = 10π; A = 25π'

OpenStudy (anonymous):

okay yeah i saw 5 so i thought maybe d.. thank you

OpenStudy (jackofalltradez):

No problem :)